A 1
team1 of :ll.1 players is so be chosen from 15 1
members. 11
n how ways rcan this,
be done If
I.. One pairticular player is always incliuded.
1
Expert's answer
2021-09-07T02:48:28-0400
"\dbinom{15-1}{11-1}=\dfrac{14!}{10!(14-10)!}"

"=\dfrac{14(13)(12)(11)}{1(2)(3)(4)}=1001"

11players from 15 members can be chosen in 1001 ways, if one particular player is always included.